### v2

# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

Abstract
We introduce CADSpotting, an efficient method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with the diversity of symbols, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing each primitive with densely sampled points, described by essential attributes like coordinates and colors. Building upon a unified 3D point cloud model for joint semantic, instance, and panoptic segmentation, CADSpotting learns robust feature representations. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce a large-scale CAD dataset named LS-CAD to support our experiments. Each floorplan in LS-CAD has an average coverage of 1,000 m2 (versus 100m2 in the existing dataset), providing a valuable benchmark for symbol spotting research. Experimental results on FloorPlanCAD and LS-CAD datasets demonstrate that CADSpotting outperforms existing methods, showcasing its robustness and scalability for real-world CAD applications.
1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilized over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpretation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
Despite the progress made in symbol recognition, the task of panoptic symbol spotting in CAD drawings remains under-explored. The challenges are multifaceted, arising from the vast diversity of symbol types, the need to differentiate between visually similar symbols, and the presence of overlapping elements. Additionally, variations in scale, orientation, and stylistic representation further complicate accurate identification.
Traditional methods approached symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 39] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 39], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of efficient strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of efficient methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder module for panoptic symbol spotting, effectively leveraging the detailed learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and efficient panoptic segmentation.
On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes. Each drawing follows the same fine-grained annotation standards as the FloorPlanCAD dataset, with an average coverage of over 1,000 square meters per floorplan. LS-CAD is the first dataset of its scale in this domain, part of which will be released to the research community under a license waiver, to stimulate further developments. Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
2 Related Work
Panoptic Image Segmentation.
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 33, 1, 35] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [35], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [40] further refines segmentation capabilities using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
CAD Symbol Spotting.
Early efforts [22, 23, 26, 10, 36] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marked a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Building on this, Zheng et al. [39] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [38, 34, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
3 Method
An overview of our CADSpotting method is provided in Fig.2. By integrating a dense point sampling strategy within a unified point cloud processing model, our approach efficiently extracts features for CAD primitives, producing a robust representation that supports various primitive types. Details are provided in Sec. 3.1. Then, a straightforward transformer decoder is applied to perform semantic, instance, and panoptic segmentation, with the corresponding loss function detailed in Sec.3.2. Additionally, we propose a novel Sliding Window Aggregation technique in Sec.3.3. Using a weighted voting strategy and Non-Maximum Suppression (NMS), this method allows our method to process expansive CAD drawings and perform large-scale panoptic segmentation.
3.1 Primitive Feature Learning
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint demonstrates strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to fully capture the diverse range of CAD symbols. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve both the efficiency and generalizability of primitive-level representations.
Dense Point Sampling.
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater efficiency and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector :
where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [34] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides high efficiency and robust performance in feature extraction.
Feature Pooling.
To obtain the final feature representation for CAD graphical primitives, we apply primitive mixed pooling to transform point-wise features into primitive-wise features. Specifically, we perform pooling on densely sampled points within each primitive , aggregating the point features of each primitive into a single primitive-wise feature vector:
where denotes the total number of 2D sampled points, and represents the number of primitives. Let represent a single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points :
This pooling method condenses a large set of dense point features into a more compact set of primitive features, effectively reducing computational complexity and storage requirements. At the same time, primitive-wise pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive level, aligning with our final goal of performing CAD spotting tasks.
3.2 Panoptic Symbol Spotting
We utilize a simple transformer decoder to extract semantic and instance symbol spotting information from the learned primitive features. In our model, the pooled primitive features serve as key and value inputs to three consecutive layers of Transformer decoders, where the self-attention mechanism effectively captures dependencies among the input features. The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
Following OneFormer3D [14], we partition the decoder output into two parts: the first outputs represent instance symbol spotting proposals, and the subsequent outputs correspond to semantic symbol spotting. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding. Here, is manually defined, and corresponds to the number of semantic categories.
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [32] to suppress non-maximal instance predictions.
Loss Function.
The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between each predicted instance and its ground truth counterpart.
3.3 Sliding Window Aggregation
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance symbol segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
Therefore, we propose using a fixed-size window based on the training image dimensions and apply 2D sliding window inference on the large input drawings. For each sliding window, once the primitives intersects with or are encompassed by the window, we utilize the features of these primitives along with the window’s semantic and instance proposals, thereby increasing the likelihood of fully capturing potential instances. After gathering proposals across all windows, we apply distinct aggregation strategies for semantic and instance proposals.
For semantic proposals, we use a voting scheme to predict each primitive’s semantic label, treating each primitive as an independent unit. Primitives that extend across larger spatial areas may be captured by multiple windows. In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
For instance proposals, we directly apply Non-Maximum Suppression (NMS) to the instance proposals obtained from all windows. A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.
| Building | Area() | #Primitive | #Instance |
| Office 1 | 3,912 | 3,910 | 406 |
| Office 2 | 1,799 | 28,193 | 1,185 |
| Office 3 | 2,994 | 16,803 | 559 |
| Campus 1 | 1,949 | 12,571 | 603 |
| Hotel 1 | 1,193 | 26,709 | 335 |
4 Experiments
4.1 LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases five large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, stairs, elevators, different types of furniture, and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
4.2 Experimental Settings
Implementation Details.
We use the train split of the FloorPlanCAD dataset to train our model. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. For dense point sampling, we set a fixed distance . Model training is performed on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of Sliding Window Aggregation on LS-CAD dataset.
Metrics.
Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ, in particular, is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
4.3 Quantitative Evaluation
| Methods | AP50 | AP75 | mAP |
| DINO [37] | 64.0 | 54.9 | 47.5 |
| SymPoint [16] | 66.3 | 55.7 | 52.8 |
| SymPoint-V2 [17] | 71.3 | 60.7 | 60.1 |
| CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
| Method | Total | Thing | Stuff | ||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| PanCADNet [5] | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
| CADTransformer [6] | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
| GAT-CADNet [39] | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
| SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
| SymPoint-V2 [17] | 90.1 | 96.3 | 93.6 | 90.8 | 96.6 | 94.0 | 80.8 | 90.9 | 88.9 |
| CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |
To demonstrate the effectiveness of our approach, we compare it against SOTA and concurrent methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
Semantic Symbol Spotting.
As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [39], SymPoint [16] and the concurrent SymPoint-V2 [17]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% and SymPoint-V2 by 8.6% in mIoU.
Instance Symbol Spotting.
We also evaluate our method against DINO [37], SymPoint and SymPoint-V2 for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As illustrated in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 0.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint-V2.
Panoptic Symbol Spotting.
We then compare our method with the same SOTA and concurrent methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Our method achieves results comparable to the concurrent SymPoint-V2, with only a slight performance gap of no more than 1.2% , except for RQ in the Stuff type, where our method outperforms. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
We also compare CADSpotting with SymPoint on our LS-CAD dataset. Since the original SymPoint cannot directly process large-scale CAD drawings, we first apply straightforward block partitioning, using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks for input into SymPoint, referred to as SymPointBP. The results in Table.7 show that our method with block partitioning (CADSpottingBP) achieves the highest performance in SQ. Our method with the SWA technique (CADSpottingSWA) achieves results comparable to SymPointBP and CADSpottingBP, with an absolute improvement of 2.5% in SQ compared to SymPointBP and a slight decrease of 0.8% compared to CADSpottingBP. However, CADSpottingSWA significantly outperforms SymPointBP in PQ and RQ, with an improvement of at least 20%, demonstrating that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
| Primitive Pooling Type | PQ | SQ | RQ |
| w/o Pooling | 74.6 | 89.3 | 83.6 |
| Max | 86.1 | 94.1 | 91.5 |
| Average | 87.6 | 94.5 | 92.7 |
| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
4.4 Qualitative Comparison
We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset. Fig.4 and Fig.5 show that our method achieves accurate and robust semantic and panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols. This demonstrates the effectiveness of our proposed dense point sampling based primitive features. Additionally, Fig.6 presents our semantic and panoptic symbol spotting results on a large-scale CAD drawing of an office building with 28,193 primitives and 1,185 instances from our LS-CAD dataset. These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
4.5 Ablation Studies
Feature Learning.
We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [38] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [34], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
| Methods | Backbone | PQ | SQ | RQ |
| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| SymPoint + Block Partitioning | 33.5 | 84.6 | 39.6 |
| Ours + Block Partitioning | 50.1 | 87.9 | 56.6 |
| Ours + SWA | 54.3 | 87.1 | 61.8 |
Pooling.
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table 5, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
Partitioning.
Table.7 shows that our method with the SWA technique (CADSpottingSWA) yields results comparable to our method using block partitioning (CADSpottingBP), with only a slight decrease of 0.8% in SQ. However, CADSpottingSWA surpasses CADSpottingBP in PQ and RQ by a margin, highlighting the effectiveness of the Sliding Window Aggregation technique in panoptic symbol spotting for large-scale CAD drawings.
5 Conclusion
We have introduced CADSpotting, a novel approach for panoptic symbol spotting in large-scale CAD drawings. Our method utilizes dense point sampling within a unified point cloud processing framework to effectively learn robust feature representations of CAD primitives. To address the challenges of precise segmentation in complex, large-scale floorplans, we propose a Sliding Window Aggregation technique, which combines weighted voting with Non-Maximum Suppression to achieve precise symbol detection and segmentation. Additionally, we present LS-CAD, a new large-scale dataset containing 50 annotated CAD floorplans from diverse building types, providing a valuable benchmark for the CAD symbol spotting task. Experimental results on both the LS-CAD and FloorPlanCAD datasets demonstrate the effectiveness, scalability, and robustness of our approach.
Despite these advancements, one limitation of our method is its panoptic symbol spotting performance on large-scale CAD drawings, which currently lags behind its performance on smaller-scale datasets like FloorPlanCAD. Addressing this gap is a focus of our ongoing and future work. We also plan to extend our approach to joint symbol spotting and 3D model generation across multiple building layers, moving towards the goal of fully automated 3D building reconstruction from CAD drawings.
Supplementary Material
Appendix A LS-CAD Dataset
In Figure 8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset. The top panel illustrates Office 3, as described in the main text, which spans a floor area of 2,994 and includes 16,803 primitives and 559 instances. The bottom panel depicts Hotel 1, also described in the main text, which spans a floor area of 1,193 and includes 26,709 primitives and 335 instances.
Appendix B Quantitative Evaluation
Detailed Ablation Study Results
We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods. Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes. The detailed results are shown in Table 9.
Appendix C Qualitative Evaluation
We provide visual comparisons of prediction results on different datasets. For FloorplanCAD, we compare the ground truth (GT) and predicted results using their primitive colors to intuitively evaluate prediction accuracy, as shown in Figure 9 and 10. For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Figure 11 and 12.
Appendix D Inference Time
We conducted a test and comparison of inference speeds for different methods on the FloorplanCAD test set using a single A6000 GPU. The results are presented in Table 8.
| Method | Backbone | Pooling | Speed |
| Handcrafted Features | PTv1 | 83ms | |
| Handcrafted Features | PTv3 | 60ms | |
| Ours | PTv3 | 133ms | |
| Ours | PTv3 | 90ms |
Appendix E Slding Window Implementation Details
In the matrix-NMS [32] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows. Figure 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation. By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
| class | A | B | C | D | ||||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| single door | 91.59 | 97.08 | 94.34 | 75.67 | 90.68 | 83.44 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
| double door | 93.51 | 96.93 | 96.47 | 75.14 | 88.76 | 84.65 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
| sliding door | 96.54 | 98.78 | 97.73 | 83.55 | 50.31 | 83.55 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
| folding door | 57.67 | 87.88 | 65.62 | 39.66 | 95.83 | 41.38 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
| revolving door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| rolling door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| window | 89.60 | 97.24 | 92.14 | 75.49 | 91.87 | 82.17 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
| bay window | 48.15 | 95.26 | 50.55 | 41.92 | 88.51 | 47.37 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
| blind window | 90.57 | 98.17 | 92.25 | 77.71 | 96.14 | 80.83 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
| opening symbol | 46.64 | 76.82 | 60.71 | 37.97 | 72.60 | 52.30 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
| sofa | 88.53 | 96.29 | 91.93 | 71.82 | 82.05 | 87.53 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
| bed | 89.79 | 92.47 | 97.11 | 78.30 | 83.76 | 93.47 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
| chair | 83.58 | 94.04 | 88.88 | 70.05 | 85.54 | 81.89 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
| table | 75.39 | 88.46 | 85.22 | 53.18 | 78.36 | 67.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
| TV cabinet | 95.23 | 97.57 | 97.60 | 82.85 | 86.79 | 95.46 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
| Wardrobe | 93.76 | 97.23 | 96.43 | 84.53 | 89.58 | 94.36 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
| cabinet | 81.40 | 91.16 | 89.28 | 65.83 | 81.45 | 80.82 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
| gas stove | 97.34 | 99.17 | 98.16 | 90.08 | 96.15 | 93.69 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
| sink | 87.59 | 95.21 | 92.00 | 76.13 | 91.10 | 83.57 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
| refrigerator | 91.96 | 96.24 | 95.55 | 71.86 | 81.87 | 87.78 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
| air conditioner | 86.53 | 99.02 | 87.39 | 78.66 | 96.36 | 81.64 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
| bath | 79.30 | 93.31 | 84.98 | 63.84 | 83.21 | 76.73 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
| bath tub | 85.00 | 91.92 | 92.47 | 70.96 | 82.90 | 85.61 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
| washing machine | 89.17 | 97.74 | 91.23 | 78.29 | 90.75 | 86.27 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
| squat toilet | 95.35 | 98.25 | 97.05 | 78.70 | 94.34 | 83.42 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
| urinal | 94.32 | 98.76 | 95.51 | 80.33 | 92.08 | 87.24 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
| toilet | 95.21 | 98.08 | 97.07 | 81.39 | 91.92 | 88.55 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
| stairs | 86.47 | 94.28 | 91.72 | 74.68 | 89.68 | 83.27 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
| elevator | 96.15 | 97.53 | 98.58 | 79.08 | 94.76 | 83.44 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
| escalator | 73.73 | 88.37 | 83.44 | 67.98 | 79.15 | 85.88 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
| row chairs | 85.16 | 94.94 | 89.71 | 82.45 | 92.43 | 89.21 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
| parking spot | 89.68 | 95.20 | 94.20 | 84.34 | 92.67 | 91.01 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
| wall | 82.37 | 88.83 | 92.73 | 75.08 | 81.61 | 66.81 | 81.86 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
| curtain wall | 64.80 | 89.88 | 72.09 | 64.17 | 84.45 | 75.98 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
| railing | 75.67 | 92.22 | 82.06 | 62.87 | 85.69 | 73.37 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
| total | 88.91 | 95.65 | 92.96 | 74.59 | 89.26 | 83.56 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |
A: Dense point sampling + PTv3 + Pooling(Ours). B: Dense point sampling + PTv3. C: Handcrafted features + PTv3. D: Handcrafted features + PTv1