### v3

# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

Abstract
We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with symbol diversity, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing primitives through densely sampled points with attributes like coordinates and colors, using a unified 3D point cloud model for robust feature learning. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce LS-CAD, a new large-scale CAD dataset to support our experiments, with each floorplan covering around 1,000 m2, significantly larger than previous benchmarks. Experiments on FloorPlanCAD and LS-CAD datasets show that CADSpotting significantly outperforms existing methods. We also demonstrate its practical value through automating parametric 3D reconstruction, enabling interior modeling directly from raw CAD inputs.
1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilize over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
Despite the progress made in symbol recognition, the task of panoptic symbol spotting in CAD drawings remains under-explored. The challenges are multifaceted, arising from the vast diversity of symbol types, the need to differentiate between visually similar symbols, and the presence of overlapping elements. Additionally, variations in scale, orientation, and stylistic representation further complicate accurate identification.
Traditional methods approach symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 40] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 40], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of effective strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of feasible methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder for panoptic symbol spotting, effectively leveraging the learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and effective panoptic segmentation.
On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes. Each drawing follows the same fine-grained annotation standards as the FloorPlanCAD dataset, with an average coverage of over 1,000 square meters per floorplan. LS-CAD is the first dataset of its scale in this domain, part of which will be released to the research community under a license waiver, to stimulate further developments. Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
We also demonstrate the effectiveness of CADSpotting by automating the generation of 3D interior models through parametric modeling. Using precise instance-level data obtained from CADSpotting, we compute spatial parameters for architectural elements, including wall contours, door and window positions, door orientations, and pivot points. With these spatial parameters clearly defined, we leverage parametric reconstruction techniques within Blender to efficiently build 3D interior models directly from raw CAD data. This approach significantly streamlines the modeling process, enhancing productivity and accuracy. To the best of our knowledge, our proposed pipeline is the first comprehensive solution that seamlessly integrates CAD segmentation with automated 3D interior model reconstruction.
2 Related Work
Panoptic Image Segmentation.
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 34, 1, 36] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [36], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [41] further refines segmentation using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
CAD Symbol Spotting.
Early efforts [22, 23, 26, 10, 37] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marks a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Countable object instances (e.g., windows, doors) can be clearly identified and individually counted whereas uncountable stuff (e.g., wall, railing) refers to elements that cannot be individually distinguished. Building on this, Zheng et al. [40] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [39, 35, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
3 Method
An overview of our CADSpotting method is illustrated in Fig.2. By incorporating a dense point sampling strategy within a unified point cloud processing model, our method effectively extracts features for CAD primitives, yielding a robust representation that supports various primitive types. We describe each module in detail in the following sections.
3.1 Primitive Feature Learning
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint shows strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to capture the diverse range of CAD symbols. Real-world CAD drawings often contain various complex elements, such as Bézier curves, rather than simple predefined line types. Therefore, methods that rely on handcrafted features lack sufficient robustness. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve generalizability of primitive-wise representations.
Dense Point Sampling.
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater feasibility and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector: ,where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [35] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides robust performance in feature extraction.
Feature Pooling.
To obtain the final feature representation for CAD primitives, we apply Primitive Mixed Pooling (PMP) to transform point-wise features into primitive-wise features. Let denote our pooling operator, which aggregates the point features into primitive-wise feature vector :
where denotes the total number of sampled points, and represents the number of primitives. Let represent the single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points:
This pooling method condenses a large set of dense point features into a more compact set of primitive-level features, effectively reducing computational complexity and storage requirements. At the same time, primitive-level pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive-level, aligning with our final goal of performing CAD spotting tasks.
3.2 Panoptic Symbol Spotting
We utilize a simple transformer decoder to extract semantic and instance symbol spotting information from the learned primitive features. In our model, the pooled primitive features serve as key and value inputs to three consecutive layers of Transformer decoders, where the self-attention mechanism effectively captures dependencies among the input features. The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
Following OneFormer3D [14], we partition the decoder output into two parts: the first represent instance proposals, and the subsequent outputs correspond to semantic symbol spotting. The output dimension of the + vector is , integrating both instance and semantic information. The first dimensions are used to generate instance masks, while the remaining dimensions determine the semantic category of each primitive. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding.
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. Additionally, we employ a bipartite matching strategy based on the Hungarian algorithm to find the optimal correspondence between predicted results and ground truth labels while training. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [33] to suppress non-maximal instance predictions.
Loss Function. The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between predicted instances and ground truth.
3.3 Sliding Window Aggregation
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
Therefore, we propose using a fixed-size window based on the training image dimensions and apply 2D sliding window inference on the large input drawings. For each sliding window, once the primitives intersects with or are encompassed by the window, we utilize the features of these primitives along with the window’s semantic and instance proposals, thereby increasing the likelihood of fully capturing potential instances. After gathering proposals across all windows, we apply distinct aggregation strategies for semantic and instance proposals.
For semantic proposals, we use a voting scheme to predict each primitive’s semantic label, treating each primitive as an independent unit. Primitives that extend across larger spatial areas may be captured by multiple windows. In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
For instance proposals, we utilize Sparse Non-Maximum Suppression(Sparse-NMS) on the proposals generated from all windows, which increases processing efficiency (detailed analysis is provided in the supplementary material). A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.
3.4 Automated 3D Interior Reconstruction
CADSpotting delivers precise panoptic segmentation results, serving as the basis for our parametric 3D interior modeling algorithm. Utilizing detailed instance information and primitive positions from segmentation, we apply instance-level processing techniques to compute spatial parameters for key architectural elements, specifically walls, doors, and windows. For doors, we determine their positions, orientations, and pivot points by analyzing the relationships between arcs and lines within each segmented instance. Window positions, on the other hand, are directly extracted from their corresponding instances. Complex wall structures require a specialized approach: we first traverse and merge endpoints of adjacent lines within wall instances, forming closed polygons. These polygons are then rasterized into binary images, from which final wall contours are identified using the method from [32]. With these spatial parameters, we reconstruct a complete 3D interior model using a parametric reconstruction pipeline in Blender. The necessary 3D primitives for walls, doors, and windows are manually prepared in Blender, providing reusable templates for streamlined parametric interior modeling. This approach significantly enhances efficiency, automating the conversion of raw CAD data into structurally accurate 3D interior models within minutes through procedural generation. An example result from our automated 3D interior reconstruction method is illustrated in Fig. 4, with additional details provided in the supplementary material.
| Building | Area() | #Primitive | #Instance |
| Office 1 | 1,799 | 28,193 | 1,185 |
| Campus 1 | 1,949 | 12,571 | 603 |
| Hotel 1 | 1,193 | 26,709 | 335 |
4 Experiments
4.1 The Proposed LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases 3 large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, elevators and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
4.2 Experimental Settings
Implementation Details. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. To ensure fair comparison, we utilize the same training subset from the FloorPlanCAD dataset as used by SymPoint. We train our model on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. For dense point sampling, we set a fixed distance . We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of SWA on LS-CAD dataset.
Metrics. Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
4.3 Quantitative Evaluation
| Methods | AP50 | AP75 | mAP |
| DINO [38] | 64.0 | 54.9 | 47.5 |
| SymPoint [16] | 66.3 | 55.7 | 52.8 |
| CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
| Method | Total | Thing | Stuff | ||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| PanCADNet [5] | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
| CADTransformer [6] | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
| GAT-CADNet [40] | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
| SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
| CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |
To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
Semantic Symbol Spotting. As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [40], and SymPoint [16]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% in mIoU.
Instance Symbol Spotting. We also evaluate our method against DINO [38] and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As shown in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 5.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint.
Panoptic Symbol Spotting. We then compare our method with the same methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
We also compare the effectiveness of Sliding Window Aggregation (SWA) and Block Partitioning (BP) methods when applied to the CADSpotting model on the LS-CAD dataset. We apply BP method using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks. When employing the BP method, we ensure that each primitive is uniquely represented by retaining only those whose starting points fall within the block. The results in Table 5 indicate that, compared to the BP method, SWA achieves the highest performance on PQ, with an absolute improvement of 10.4%, although there is a slight decrease of 1.0% on SQ. The result shows that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
4.4 Qualitative Comparison
We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset. Fig.5 shows that our method achieves accurate and robust panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols. This demonstrates the effectiveness of our proposed dense point sampling based primitive features. Additionally, Fig.6 presents our semantic and panoptic symbol spotting results on a large-scale CAD drawing of an office building with 28,193 primitives and 1,185 instances from our LS-CAD dataset. These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
| Method | PQ | SQ | RQ |
| Ours + BP | 42.8 | 94.8 | 45.2 |
| Ours + SWA | 53.2 | 93.8 | 56.6 |
To evaluate model generalizability and demonstrate the utility of LS-CAD dataset, we also train models on a combined dataset consisting of FloorPlanCAD and LS-CAD. Comprehensive experimental results and analyses are provided in the supplementary material.
| Methods | Backbone | PQ | SQ | RQ |
| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| w/o Color Feature | 85.8 | 92.5 | 92.7 |
| Color Feature | 88.9 | 95.6 | 93.0 |
4.5 Ablation Studies
In this section, we present ablation studies examining the effects of feature learning and RGB information. Additional ablation studies, which explore various pooling strategies and evaluate performance across all classes, are provided in the supplementary material.
Feature Learning. We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [39] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [35], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
RGB Color Information. We compare the performance of CADSpotting through the inclusion versus exclusion of color information in the FloorPlanCAD dataset. Table.7 shows that adding RGB information improves the PQ metric by 3.1%.
5 Conclusion
We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings. Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation. We also present LS-CAD, a dataset of 50 annotated floorplans from diverse building types. Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
Our approach currently faces limitations due to the dataset’s primary emphasis on residential and commercial buildings, which restricts its applicability to industrial and cultural contexts. Additionally,it may struggle to capture significant stylistic variations among CAD symbols. Addressing these challenges through the integration of in-context learning may help enhance annotation efficiency and extend applicability. This is left as our future work.
Supplementary Material
| Method | Training Dataset | Test on FloorPlanCAD | Test on Cross-dataset | ||||
| PQ | SQ | RQ | PQ | SQ | RQ | ||
| SymPoint [16] | FloorPlanCAD | 83.3 | 91.4 | 91.1 | 33.2 | 83.0 | 39.9 |
| SymPoint | Cross-dataset | 79.0 | 89.2 | 88.6 | 57.0 | 84.0 | 67.8 |
| CADSpotting | FloorPlanCAD | 88.9 | 95.6 | 93.0 | 42.8 | 94.8 | 45.2 |
| CADSpotting | Cross-dataset | 87.3 | 95.4 | 91.4 | 69.9 | 91.6 | 76.3 |
| CADSpotting(SWA) | FloorPlanCAD | / | 53.2 | 93.8 | 56.6 | ||
| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 84.9 | 88.9 | ||
6 More Details on the LS-CAD Dataset
In Fig.8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset. The top panel illustrates Office 3, which spans a floor area of 2,994, including 16,803 primitives and 559 instances. The bottom panel depicts Hotel 1, which spans a floor area of 1,193, consisting of 26,709 primitives and 335 instances.
To demonstrate the generalization capability of our model and highlight the value of the LS-CAD dataset, we implement a cross-dataset joint training paradigm. We randomly partition complete large-scale CAD drawings into three subsets: 80% for training, 10% for validation, and 10% for testing. The division of the LS-CAD dataset strictly adheres to the standardized data splitting protocol established in FloorPlanCAD. The training and validation sets from both datasets are subsequently integrated to construct a cross-dataset collaborative training environment. As shown in Table.8, the cross-dataset joint training strategy significantly improves model performanc on the LS-CAD test set, although with slight performance degradation on the original FloorPlanCAD test set. This demonstrates the necessity of the proposed LS-CAD dataset for provding more diverse CAD representations. Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric. After cross-dataset joint training, the PQ score for CADSpotting decreases by only 1.2% on FloorPlanCAD, whereas it drops by 4.3% on SymPoint. This indicates that our model exhibits stronger generalization capabilities.
7 Quantitative Evaluation
Ablation Study on Feature Learning.
We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods. Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes. The detailed results are shown in Table.10.
Ablation Study on Pooling.
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table.9, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
8 Qualitative Evaluation
We provide comparisons of prediction results on different datasets. For FloorPlanCAD, we compare the ground truth and prediction using their primitive colors to intuitively evaluate prediction accuracy, as shown in Fig.9, Fig.10 and Fig.13. For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Fig.11 and Fig.12.
9 Sliding Window Implementation Details
The Algorithm.1 details the implementation process of our SWA. In the matrix-NMS [33] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows. Fig. 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation. By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
10 Automated 3D Interior Reconstruction
Based on the spotting results generated by CADSpotting, we present more 3D reconstruction renderings in Fig.14. By systematically optimizing the parameters of predefined instance components using instance segmentation data and primitive spatial coordinates, we achieve accurate 3D model reconstruction through component-level geometric adaptation
For wall polygon extraction, we employ an innovative SVG-PNG conversion strategy to differentiate floor areas from walls: 1) Merge adjacent endpoints and remove duplicate segments to simplify SVG paths. 2) Convert vectors to 8K resolution binarized PNG while maintaining coordinate alignment. 3) Identify the largest connected component as the floor through connected component analysis, with secondary components as candidate walls. 4) Re-vectorize boundaries through coordinate mapping to generate closed wall polygons.
The proposed method exhibits notable robustness against recognition inaccuracies – incomplete walls are automatically reconciled during stage 1, while small noise particles are filtered in stage 3 using area thresholds.
For doors and windows, we employ category-specific parameterization strategies. Doors are systematically categorized into four distinct subtypes (single/double/sliding/folding) based on semantic labels, with positions, orientations, and pivot points determined through linear-arc spatial relationships.
Windows are grouped via union-find algorithms, calculating average lines of co-grouped segments as representative lines, then selecting the segment closest to the representative line within each group as the window instance position.
The final reconstruction leverages Blender’s geometry nodes for procedural modeling: 1) Extrude wall polygons to 3D volumes using floorplan height attributes. 2) Instance-place door/window assemblies at infered positions. 3) Apply floor based on origin CAD drawing.
This integrated approach achieves automated 3D reconstruction with exceptional efficiency, transforming raw CAD data into structurally complete 3D interior models within minutes through streamlined procedural generation.
| Primitive Pooling Type | PQ | SQ | RQ |
| w/o Pooling | 74.6 | 89.3 | 83.6 |
| Max | 86.1 | 94.1 | 91.5 |
| Average | 87.6 | 94.5 | 92.7 |
| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
| class | A | B | C | D | ||||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| single door | 91.59 | 97.08 | 94.34 | 75.67 | 90.68 | 83.44 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
| double door | 93.51 | 96.93 | 96.47 | 75.14 | 88.76 | 84.65 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
| sliding door | 96.54 | 98.78 | 97.73 | 83.55 | 50.31 | 83.55 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
| folding door | 57.67 | 87.88 | 65.62 | 39.66 | 95.83 | 41.38 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
| revolving door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| rolling door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| window | 89.60 | 97.24 | 92.14 | 75.49 | 91.87 | 82.17 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
| bay window | 48.15 | 95.26 | 50.55 | 41.92 | 88.51 | 47.37 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
| blind window | 90.57 | 98.17 | 92.25 | 77.71 | 96.14 | 80.83 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
| opening symbol | 46.64 | 76.82 | 60.71 | 37.97 | 72.60 | 52.30 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
| sofa | 88.53 | 96.29 | 91.93 | 71.82 | 82.05 | 87.53 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
| bed | 89.79 | 92.47 | 97.11 | 78.30 | 83.76 | 93.47 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
| chair | 83.58 | 94.04 | 88.88 | 70.05 | 85.54 | 81.89 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
| table | 75.39 | 88.46 | 85.22 | 53.18 | 78.36 | 67.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
| TV cabinet | 95.23 | 97.57 | 97.60 | 82.85 | 86.79 | 95.46 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
| Wardrobe | 93.76 | 97.23 | 96.43 | 84.53 | 89.58 | 94.36 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
| cabinet | 81.40 | 91.16 | 89.28 | 65.83 | 81.45 | 80.82 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
| gas stove | 97.34 | 99.17 | 98.16 | 90.08 | 96.15 | 93.69 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
| sink | 87.59 | 95.21 | 92.00 | 76.13 | 91.10 | 83.57 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
| refrigerator | 91.96 | 96.24 | 95.55 | 71.86 | 81.87 | 87.78 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
| air conditioner | 86.53 | 99.02 | 87.39 | 78.66 | 96.36 | 81.64 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
| bath | 79.30 | 93.31 | 84.98 | 63.84 | 83.21 | 76.73 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
| bath tub | 85.00 | 91.92 | 92.47 | 70.96 | 82.90 | 85.61 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
| washing machine | 89.17 | 97.74 | 91.23 | 78.29 | 90.75 | 86.27 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
| squat toilet | 95.35 | 98.25 | 97.05 | 78.70 | 94.34 | 83.42 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
| urinal | 94.32 | 98.76 | 95.51 | 80.33 | 92.08 | 87.24 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
| toilet | 95.21 | 98.08 | 97.07 | 81.39 | 91.92 | 88.55 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
| stairs | 86.47 | 94.28 | 91.72 | 74.68 | 89.68 | 83.27 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
| elevator | 96.15 | 97.53 | 98.58 | 79.08 | 94.76 | 83.44 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
| escalator | 73.73 | 88.37 | 83.44 | 67.98 | 79.15 | 85.88 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
| row chairs | 85.16 | 94.94 | 89.71 | 82.45 | 92.43 | 89.21 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
| parking spot | 89.68 | 95.20 | 94.20 | 84.34 | 92.67 | 91.01 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
| wall | 82.37 | 88.83 | 92.73 | 75.08 | 81.61 | 66.81 | 81.86 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
| curtain wall | 64.80 | 89.88 | 72.09 | 64.17 | 84.45 | 75.98 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
| railing | 75.67 | 92.22 | 82.06 | 62.87 | 85.69 | 73.37 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
| total | 88.91 | 95.65 | 92.96 | 74.59 | 89.26 | 83.56 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |
A: Dense point sampling + PTv3 + Pooling(Ours). B: Dense point sampling + PTv3. C: Handcrafted features + PTv3. D: Handcrafted features + PTv1